A 20.0 cm × 20.0 cm square loop of wire lies in the xy-plane with its bottom edge on the x-axis. The resistance of the loop is 0.530 Ω . A magnetic field parallel to the z-axis is given by B = 0.850 y2t, where B is in tesla, y is in meters, and t is in seconds. What is the size of the induced current in the loop at t = 0.460 s ?
Answer #1

Magnetic flux through the loop is

Induced emf

E=0.85*0.2*0.23/3

E=4.533*10-4 V

Induced current is

I=E/R =(4.533*10-4)/0.53

I=8.553*10-4 A or 0.8553 mA